A single-turn square loop 20.0 cm on a side carries a 1.5 A current.


(a) What’s the loop’s magnetic dipole moment?


(b) What’s the magnitude of the torque the loop would experience if it was in a 2.3T magnetic field with the loop’s dipole moment vector at 60^◦ to the field?


1
Expert's answer
2021-03-09T07:42:15-0500

a)

"\mu=iA"


"i(l\times w)"


"1.5\times(20\times20)"


"= 600 A.m^{2}"


b)

"T=\mu Bsin(\theta)"

"=600\times2.3\times sin(60)"

"=1195.12N.m"